OpenStudy (anonymous):

how do I find were P converges in the series of: the sum from n=1 to infinity of (7n^7/(4n^5+2))^p

OpenStudy (anonymous):

the degree of the denominator has to be more than one larger than the degree of the numerator

OpenStudy (anonymous):

for example 2 would work because then the degree of the denominator would be 10 and that is 3 larger than 7 of course you can pick a smaller p

OpenStudy (anonymous):

ca you explain a little more? i haven't fully grasped the concept

ganeshie8 (ganeshie8):

\[\sum\limits_{n=1}^{\infty}\left(\dfrac{7n^7}{4n^5+2}\right)^p\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need to find all values were p converges

ganeshie8 (ganeshie8):

so that exponent p is not just for the denominator term its for the entire fraction is it ?

OpenStudy (anonymous):

yes it is for the entire fraction

OpenStudy (anonymous):

\[\sum\limits_{n=1}^{\infty}\left(\dfrac{7n^7}{4n^5+2}\right)^p\] The series will diverge if the numerator degree is \(k\ge1\). If we set \(p=1\), then we can see that the series diverges because we have a "cumulative" degree of \(k=2\). (In other words, dividing a 7th degree polynomial by a 5th degree polynomial yields a 2nd degree polynomial.) You want to find \(p\) such that you can essentially normalize this degree so that it always is less than 1. Hint: Think rational numbers. You can ignore most of the algebra involved with expansions. What value of \(p\) times 2 gives you something less than 1?

ganeshie8 (ganeshie8):

you may use p-series test here just notice that "highest exponent" of numerator = 7p "highest exponent" of denominator = 5p

ganeshie8 (ganeshie8):

\[\sum\limits_{n=1}^{\infty}\left(\dfrac{7n^7}{4n^5+2}\right)^p \lt C\sum\limits_{n=1}^{\infty}\dfrac{1}{n^{-2p}}\] appealing to p-series test gives you the same answer as siths

OpenStudy (anonymous):

oh okay I see